Obtain an equation for an ellipse with major axis of lenghth 12 and foci at (-2, 5) and (-6, 5).


um i've tried solving it and at best, i've gotten the start with:


(x+4)^2 + (y-5)^2
------- -------
36 22 (?)


and there is no multiple choice answer with
36y^2,


for the y^2, the choices are......
52y^2
8y^2
9y^2


please help me soon ! :)
i need this ready by tomorrow afternoonish thank you.

Answer by scott8148(6628)

the distance between the foci is 4, so c=2

b^2=a^2-c^2 __ b^2=36-4 __ b^2=32 (not 22)

[(x+4)^2]/36+[(y-5)^2]/32=1 __ multiplying by 36*8 (LCM) __ 8(x+4)^2+9(y-5)^2=36*8